# DepthSplat: Connecting Gaussian Splatting and Depth

### v1

@@ Abstract
Gaussian splatting and single/multi-view depth estimation are typically studied in isolation. In this paper, we present DepthSplat to connect Gaussian splatting and depth estimation and study their interactions. More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions. We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabelled datasets. We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments. Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks. Our code, models, and video results are available at https://haofeixu.github.io/depthsplat/.
@@ 1 Introduction
Novel view synthesis (Buehler et al., 2001; Zhou et al., 2018) and depth prediction (Schönberger et al., 2016; Eigen et al., 2014) are two fundamental tasks in computer vision, serving as the driving force behind numerous applications ranging from augmented reality to robotics and autonomous driving. There have been notable advancements in both areas recently.
For novel view synthesis, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a popular technique due to its impressive real-time performance while attaining high visual fidelity. Recently, advances in feed-forward 3DGS models (Charatan et al., 2024; Chen et al., 2024; Szymanowicz et al., 2024) have been introduced to alleviate the need for tedious per-scene optimization, also enabling few-view 3D reconstruction. The state-of-the-art sparse-view method MVSplat (Chen et al., 2024) relies on feature matching-based multi-view depth estimation (Xu et al., 2023) to localize the 3D Gaussian positions, which makes it suffer from similar limitations (e.g., occlusions, texture-less regions, and reflective surfaces) as other multi-view depth methods (Schönberger et al., 2016; Yao et al., 2018; Gu et al., 2020; Wang et al., 2021; Duzceker et al., 2021).
On the other hand, significant progress has been made in monocular depth estimation, with recent models (Yang et al., 2024a; Ke et al., 2024; Yin et al., 2023; Fu et al., 2024; Ranftl et al., 2020; Eftekhar et al., 2021) achieving robust predictions on diverse in-the-wild data. However, these depths typically lack consistency across views, constraining their performance in downstream tasks (Wang et al., 2023; Yin et al., 2022). In addition, both state-of-the-art multi-view (Cao et al., 2022; Xu et al., 2023) and monocular (Yang et al., 2024a; Piccinelli et al., 2024; Yang et al., 2024b) depth models are trained with ground truth depth supervision, which prevents exploiting large unlabelled datasets for more robust depth predictions.
The integration of 3DGS with single/multi-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths. To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust single/multi-view depth estimation to improve the performance for both tasks.
Specifically, we first contribute a robust multi-view depth model by integrating pre-trained monocular depth features (Yang et al., 2024b) to the multi-view feature matching branch, which not only maintains the consistency of multi-view depth models but also leads to more robust results in situations that are hard to match (e.g., occlusions, texture-less regions and reflective surfaces). The predicted multi-view depth maps are then unprojected to 3D as the Gaussian centers, and we use an additional lightweight network to predict other remaining Gaussian parameters. They are combined together to achieve novel view synthesis with the splatting operation (Kerbl et al., 2023).
@@
We conduct extensive experiments on the large-scale TartanAir (Wang et al., 2020), ScanNet (Dai et al., 2017) and RealEstate10K (Zhou et al., 2018) datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV (Ling et al., 2023) dataset, which features complex real-world scenes and thus is more challenging. Under various evaluation settings, our DepthSplat achieves state-of-the-art results. The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth estimation.
@@ 2 Related Work
Multi-View Depth Estimation. As a core component of classical multi-view stereo pipelines (Schönberger et al., 2016), multi-view depth estimation exploits multi-view photometric consistency across multiple images to perform feature matching and predict the depth map of the reference image. Recently, many learning-based methods (Yao et al., 2018; Gu et al., 2020; Wang et al., 2021; Duzceker et al., 2021; Wang et al., 2022; Ding et al., 2022; Cao et al., 2022) have been proposed to improve depth accuracy. For example, MVSNet (Yao et al., 2018) uses the plane-sweep algorithm (Collins, 1996) to build a 3D cost volume, regularizes it with a 3D CNN, and then regresses the depth map. Though these learning-based methods significantly improve the depth quality when compared to traditional methods (Galliani et al., 2015; Schönberger et al., 2016; Xu & Tao, 2019), they cannot handle challenging situations where the multi-view photometric consistency assumption is not guaranteed, e.g., occlusions, low-textured areas, and non-Lambertian surfaces.
@@
Monocular Depth Estimation. Recently, we have witnessed significant progress in depth estimation from a single image (Ranftl et al., 2020; Bhat et al., 2023; Yin et al., 2023; Yang et al., 2024a; Ke et al., 2024), and existing methods can produce surprisingly accurate results on diverse in-the-wild data. However, monocular depth methods inherently suffer from scale ambiguities and are not able to produce multi-view consistent depth predictions, which are crucial for downstream tasks like 3D reconstruction (Yin et al., 2022) and video depth estimation (Wang et al., 2023). In this paper, we leverage the powerful features from a pre-trained monocular depth model (Yang et al., 2024b) to augment feature-matching based multi-view depth estimation, which not only maintains high multi-view consistency but also leads to significantly improved robustness in challenging situations such as low-textured regions and reflective surfaces.
Monocular and Multi-View Depth Fusion. Several previous methods (Bae et al., 2022; Li et al., 2023; Yang et al., 2022; Cheng et al., 2024) try to fuse monocular and multi-view depths to alleviate the limitations of feature matching-based multi-view depth estimations. However, they either fuse single- and multi-view depth predictions with additional networks or rely on sophisticated architectures. In contrast, we identify the power of off-the-shell pre-trained monocular depth models and propose to augment multi-view cost volumes with monocular features, which not only leads to a simple model architecture but also strong performance.
@@
Feed-Forward Gaussian Splatting. Several feed-forward 3D Gaussian splatting models (Charatan et al., 2024; Szymanowicz et al., 2024; Chen et al., 2024; Wewer et al., 2024; Tang et al., 2024; Xu et al., 2024b; Zhang et al., 2024) have been proposed in literature thanks to its efficiency and ability to handle sparse views. In particular, pixelSplat (Charatan et al., 2024) and Splatter Image (Szymanowicz et al., 2024) predict 3D Gaussians from image features, while MVSplat (Chen et al., 2024) encodes the feature matching information with cost volumes and achieves better geometry. However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces. In this paper, we propose to integrate monocular features from pre-trained monocular depth models for more robust depth prediction and Gaussian splatting reconstruction. We additionally study the interactions between depth and Gaussian splatting tasks. Another line of work like LGM (Tang et al., 2024), GRM (Xu et al., 2024b), and GS-LRM (Zhang et al., 2024) relies significantly on the training data and compute, discarding explicit feature matching cues and learning everything purely from data. This makes them expensive to train (e.g., GS-LRM (Zhang et al., 2024) is trained with 64 A100 GPUs for 2 days), while our model can be trained in 1 day with 8 GPUs. Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabelled datasets without the need for ground truth depths.
Depth and Gaussian Splatting. In addition to the experimental setup (feed-forward) studied in this paper, recently, there has been another line of work (Chung et al., 2024; Turkulainen et al., 2024) which applies an additional depth loss in the Gaussian splatting optimization process. We remark that these two approaches (feed-forward vs. per-scene optimization) are orthogonal. In particular, our approach focuses on exploring the advanced network architectures and the power of large-scale training data, while the optimization methods mainly study the effect of loss functions for regularizing the optimization process.
@@ 3.1 Multi-View Feature Matching
Multi-View Feature Extraction. For $N$ input images, we first use a lightweight weight-sharing ResNet (He et al., 2016) architecture to get $s\times$ downsampled features for each image independently. To handle different image resolutions, we make the downsampling factor $s$ flexible by controlling the number of stride-2 $3\times 3$ convolutions. For example, the downsampling factor $s$ is 4 if two stride-2 convolutions are used and 8 if three are used. To exchange information across different views, we use a multi-view Swin Transformer (Liu et al., 2021; Xu et al., 2022; 2023) which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension. More specifically, cross-attention is performed between each reference view and other views. When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera position distances to the reference view. This makes the computation tractable while maintaining cross-view interactions.
Feature Matching. We encode the feature matching information across different views with the plane-sweep stereo approach (Collins, 1996; Xu et al., 2023). More specifically, for each view $i$, we first uniformly sample $D$ depth candidates $\{d_{m}\}_{m=1}^{D}$ from the near and far depth ranges and then warp the feature ${\bm{F}}^{j}$ of view $j$ to view $i$ with the camera projection matrix and each depth candidate $d_{m}$. Then we obtain $D$ warped features $\{{\bm{F}}^{j\to i}_{d_{m}}\}_{m=1}^{D}$ that correspond to feature ${\bm{F}}^{i}$. We then measure their feature correlations with the dot-product operation (Xu & Zhang, 2020; Chen et al., 2024). The cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ for image $i$ is obtained by stacking all correlations. Accordingly, we obtain cost volumes $\{{\bm{C}}_{i}\}_{i=1}^{N}$ for all input images $\{{\bm{I}}_{i}\}_{i=1}^{N}$. For more than two input views, similar to the strategy in cross-view attention computation, we select the top-$2$ nearest views for each reference view and compute feature correlations with only the selected views. This enables our cost volume construction to achieve a good speed-accuracy trade-off and scale efficiently to a larger number of input views. The correlation values for the two selected views are combined with averaging.
@@ 3.2 Monocular Depth Feature Extraction
Despite the remarkable progress in multi-view feature matching-based depth estimation (Yao et al., 2018; Wang et al., 2022; Xu et al., 2023) and Gaussian splatting (Chen et al., 2024), they inherently suffer from limitations in challenging situations like occlusions, texture-less regions, and reflective surfaces. Thus, we propose to integrate pre-trained monocular depth features into the cost volume representation to handle scenarios that are challenging or impossible to match.
More specifically, we leverage the pre-trained monocular depth backbone from the recent Depth Anything (Yang et al., 2024b) model thanks to its impressive performance on diverse in-the-wild data. The monocular backbone is a ViT (Dosovitskiy et al., 2020; Oquab et al., 2023) model, which has a patch size of 14 and outputs a feature map that is $1/14$ spatial resolution of the original image. We simply bilinearly interpolate the spatial resolution of the monocular features to the same resolution as the cost volume in Sec. 3.1 and obtain the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ for input image ${\bm{I}}_{i}$, where $C_{\mathrm{mono}}$ is the dimension of the monocular feature. This process is performed for all input images in parallel and we obtain monocular features $\{{\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}\}_{i=1}^{N}$, which are subsequently used for per-view depth map estimations.
@@ 3.3 Feature Fusion and Depth Regression
To achieve robust and multi-view consistent depth predictions, we combine the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ and cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ via simple concatenation in the channel dimension. A subsequent 2D U-Net (Rombach et al., 2022; Ronneberger et al., 2015) is used to regress depth from the concatenated monocular features and cost volumes. This process is performed for all the input images in parallel and for each image, it outputs a tensor of shape $\frac{H}{s}\times\frac{W}{s}\times D$, where $D$ is the number of depth candidates. We then normalize the $D$ dimension with the softmax operation and perform a weighted average of all depth candidates to obtain the depth output.
We also apply a hierarchical matching (Gu et al., 2020) architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further. More specifically, based on the coarse depth prediction, we perform a correspondence search on the 2×2$\times$2 × higher feature maps within the neighbors of the 2×2$\times$2 × upsampled coarse depth prediction. Since we already have a coarse depth prediction, we only need to search a smaller range at the higher resolution, and thus, we construct a smaller cost volume compared to the coarse stage. Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features (Gu et al., 2020). Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled to the full resolution with a learned upsampler (Ranftl et al., 2021).

### v2

@@ Abstract
Gaussian splatting and single/multi-view depth estimation are typically studied in isolation. In this paper, we present DepthSplat to connect Gaussian splatting and depth estimation and study their interactions. More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions. We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabeled datasets. We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments. Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks.
@@ 1 Introduction
Novel view synthesis [4, 75] and depth prediction [41, 18] are two fundamental tasks in computer vision, serving as the driving force behind numerous applications ranging from augmented reality to robotics and autonomous driving. There have been notable advancements in both areas recently.
For novel view synthesis, 3D Gaussian Splatting (3DGS) [27] has emerged as a popular technique due to its impressive real-time performance while attaining high visual fidelity. Recently, advances in feed-forward 3DGS models [7, 8, 46] have been introduced to alleviate the need for tedious per-scene optimization, also enabling few-view 3D reconstruction. The state-of-the-art sparse-view method MVSplat [8] relies on feature matching-based multi-view depth estimation [61] to localize the 3D Gaussian positions, which makes it suffer from similar limitations (e.g., occlusions, texture-less regions, and reflective surfaces) as other multi-view depth methods [41, 67, 23, 53, 16].
On the other hand, significant progress has been made in monocular depth estimation, with recent models [65, 26, 69, 20, 37, 17] achieving robust predictions on diverse in-the-wild data. However, these depths typically lack consistent scales across multiple views, constraining their performance in downstream tasks like 3D reconstruction [56, 68].
The integration of 3DGS with single/multi-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths. To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust monocular depth estimation to improve the performance for both tasks.
Specifically, we first contribute a robust multi-view depth model by integrating pre-trained monocular depth features [66] to the multi-view feature matching branch, which not only maintains the consistency of multi-view depth models but also leads to more robust results in situations that are hard to match (e.g., occlusions, texture-less regions and reflective surfaces). The predicted multi-view depth maps are then unprojected to 3D as the Gaussian centers, and we use an additional lightweight network to predict other remaining Gaussian parameters. They are combined together to achieve novel view synthesis with the splatting operation [27].
While previous methods [1, 29, 9] also try to fuse monocular and multi-view depths, they usually rely on sophisticated architectures. In contrast, we identify the power of off-the-shelf pre-trained monocular depth models and propose to augment multi-view cost volumes with monocular features, leading to a simpler model and stronger performance.
@@
We conduct extensive experiments on the large-scale TartanAir [55], ScanNet [12] and RealEstate10K [75] datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV [30] dataset, which features complex real-world scenes and thus is more challenging. Under various evaluation settings, our DepthSplat achieves state-of-the-art results. We also achieve promising feed-forward reconstruction results of large-scale or 360 scenes from up to 12 input views at $512\times 960$ resolutions. The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth.
@@ 2 Related Work
Multi-View Depth Estimation. As a core component of classical multi-view stereo pipelines [41], multi-view depth estimation exploits multi-view photometric consistency across multiple images to perform feature matching and predict the depth map of the reference image. In recent years, many learning-based methods [67, 23, 53, 16, 54, 13, 6] have been proposed to improve depth accuracy. Though these learning-based methods significantly improve the depth quality compared to traditional methods [21, 41, 63], they cannot handle challenging situations where the multi-view photometric consistency assumption is not guaranteed, e.g., occlusions, low-textured areas, and non-Lambertian surfaces.
@@
Monocular Depth Estimation. Recently, we have witnessed significant progress in depth estimation from a single image [37, 2, 69, 65, 26], and existing methods can produce surprisingly accurate results on diverse in-the-wild data. However, monocular depth methods inherently suffer from scale ambiguities, which makes it challenging to use them for downstream tasks like 3D reconstruction [68]. In this paper, we leverage the powerful features from a pre-trained monocular depth model [66] to augment feature-matching based multi-view depth estimation, which not only maintains the multi-view consistency but also leads to significantly improved robustness in challenging situations such as low-textured regions and reflective surfaces.
@@
Feed-Forward Gaussian Splatting. Several feed-forward 3D Gaussian splatting models [7, 46, 8, 58, 48, 64, 72] have been proposed in literature thanks to its efficiency and ability to handle sparse views. In particular, pixelSplat [7] and Splatter Image [46] predict 3D Gaussians from image features, while MVSplat [8] encodes the feature matching information with cost volumes and achieves better geometry. However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces. In this paper, we propose to integrate monocular features from pre-trained monocular depth models [66] for more robust depth prediction and 3D Gaussian reconstruction. Another line of work like LGM [48], GRM [64], and GS-LRM [72] relies significantly on the training data and compute, discarding explicit feature matching cues and learning priors purely from data. This makes them expensive to train (e.g., GS-LRM [72] is trained with 64 A100 GPUs for 2 days), while our model can be trained in 20 hours with 8 GPUs. Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabeled datasets without the need for ground truth depths.
Depth and Gaussian Splatting. Flash3D [45] explores a pre-trained monocular depth model [36] for single-image 3D Gaussian reconstruction. In contrast, we target large-scale scene reconstruction from multiple views. TranSplat [71] leverages monocular depth to improve Gaussian reconstruction. However, a crucial difference is that TranSplat uses monocular features to refine the coarse depth predicted from a cost volume, which makes it inherently vulnerable to the error in the coarse stage and, thus, suffers from the error propagation issue. In contrast, we avoid this by combining cost volume and monocular features early, which also enables our method to perform significantly better than TranSplat (Tab. 5). There has been another line of work [10, 50] which applies an additional depth loss in the Gaussian splatting optimization process. We note that these two approaches (feed-forward vs. per-scene optimization) are orthogonal.
@@ 3.1 Multi-View Feature Matching
Multi-View Feature Extraction. For $N$ input images, we first use a lightweight weight-sharing ResNet [25] architecture to get $s\times$ downsampled features for each image independently. To handle different image resolutions, we make the downsampling factor $s$ flexible by controlling the number of stride-2 $3\times 3$ convolutions. For example, the downsampling factor $s$ is 4 if two stride-2 convolutions are used and 8 if three are used. To exchange information across different views, we use a multi-view Swin Transformer [31, 60, 61] which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension. More specifically, cross-attention is performed between each reference view and other views. When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera distances to the reference view. This makes the computation tractable while maintaining cross-view interactions.
Feature Matching. We encode the feature matching information across different views with the plane-sweep stereo approach [11, 61]. More specifically, for each view $i$, we first uniformly sample $D$ depth candidates $\{d_{m}\}_{m=1}^{D}$ from the near and far depth ranges and then warp the feature ${\bm{F}}^{j}$ of view $j$ to view $i$ with the camera projection matrix and each depth candidate $d_{m}$. Then we obtain $D$ warped features $\{{\bm{F}}^{j\to i}_{d_{m}}\}_{m=1}^{D}$ that correspond to feature ${\bm{F}}^{i}$. We then measure their feature correlations with the dot-product operation [59, 8]. The cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ for image $i$ is obtained by stacking all correlations. Accordingly, we obtain cost volumes $\{{\bm{C}}_{i}\}_{i=1}^{N}$ for all input images $\{{\bm{I}}_{i}\}_{i=1}^{N}$. For more than two input views, similar to the strategy in cross-view attention computation, we select the top-$2$ nearest views for each reference view and compute feature correlations with only the selected views. This enables our cost volume construction to achieve a good speed-accuracy trade-off and scale efficiently to a larger number of input views. The correlation values for the two selected views are combined with averaging.
@@ 3.2 Monocular Depth Feature Extraction
Despite the remarkable progress in multi-view feature matching-based depth estimation [67, 54, 61] and Gaussian splatting [8], they inherently suffer from limitations in challenging situations like occlusions, texture-less regions, and reflective surfaces. Thus, we propose to integrate pre-trained monocular depth features into the cost volume to handle scenarios that are challenging or impossible to match.
More specifically, we leverage the pre-trained monocular depth backbone from the recent Depth Anything V2 [66] model thanks to its impressive performance on diverse in-the-wild data. The monocular backbone is a ViT [14, 34] model, which has a patch size of 14 and outputs a feature map that is $1/14$ spatial resolution of the original image. We simply bilinearly interpolate the spatial resolution of the monocular features to the same resolution as the cost volume in Sec. 3.1 and obtain the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ for input image ${\bm{I}}_{i}$, where $C_{\mathrm{mono}}$ is the dimension of the monocular feature. This process is performed for all input images in parallel and we obtain monocular features $\{{\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}\}_{i=1}^{N}$, which are subsequently used for per-view depth map estimations.
@@ 3.3 Feature Fusion and Depth Regression
To achieve robust and multi-view consistent depth predictions, we combine the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ and cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ via simple concatenation in the channel dimension. A subsequent 2D U-Net [39, 40] is used to regress depth from the concatenated monocular features and cost volumes. This process is performed for all the input images in parallel and for each image, it outputs a tensor of shape $\frac{H}{s}\times\frac{W}{s}\times D$, where $D$ is the number of depth candidates. We normalize the $D$ dimension with the softmax operation and perform a weighted average of all depth candidates to obtain the depth output.
We also apply a hierarchical matching [23] architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further. More specifically, based on the coarse depth prediction, we perform a correspondence search on the 2×2$\times$2 × higher feature maps within the neighbors of the 2×2$\times$2 × upsampled coarse depth prediction. Since we already have a coarse depth prediction, we only need to search a smaller range at the higher resolution, and thus, we construct a smaller cost volume compared to the coarse stage. Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features [23]. Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled [38] to the full resolution.